You have $88 altogether. The amount in the bank is 10 times as much as what's in your pocket. How much money is in each place.
Pocket........... Bank.......

Answer by addingup(3677)

bank: b
pocket:p
you can change this to x and y if that's what they are asking you to do.
:
b+p = 88
and
b = 10p
:
p+10p = 88
11p = 88
p = 8
-----------------
Check:
in your pocket: 8
in the bank 10 times more: 8*10= 80
8+80 = 88 Correct.
:
John